OpenStudy (anonymous):

1/4a=3 A.a=0.75 B.a=3 C.a=12 D.a=16 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (anonymous):

is it 1/4a or is a in the middle?

OpenStudy (anonymous):

its right next to 1/4

OpenStudy (anonymous):

a= 12

OpenStudy (anonymous):

so its C

OpenStudy (texaschic101):

1/4a = 3 divide both sides by 1/4 1/4a/(1/4) = 3/(1/4) a = 3 * 4/1 a = 12